1) Factor by grouping:
18a^3– 24a^2b + 15ab^2– 20b^3


2) Solve the following:
18x^2 + 9x – 2 = 0

3)Multiply: (x + 2y)^4

4)Simplify the following expression:
(1/2)^3 + (1/3)^-2 - (1/4)^-2

Answer by stanbon(75887)

1) Factor by grouping:
18a^3– 24a^2b + 15ab^2– 20b^3
= 6a^2(3a-4b) + 5b^2(3a-4b)
=(3a-4b)(6a^2+5b^2)
==========================


2) Solve the following:
18x^2 + 9x – 2 = 0
18x^2+12x-3x-2 = 0
6x(3x+2) - (3x+2) =0
(3x+2)(6x-1) =0
x=-2/3 or x= 1/6
==========================

3)Multiply: (x + 2y)^4
= 4C4x^4 + 4C3x^3(2y) + 4C2x^2(2y)^2 + 4C1x(2y)^3 + 4C0(2y)^4
= x^4 + 4x^3(2y) + 6x^2(4y^2) + 4x(8y^3) + 16y^4
= x^4 + 8x^3y + 24x^2y^2 + 32xy^3 + 16y^4
==================

4)Simplify the following expression:
(1/2)^3 + (1/3)^-2 - (1/4)^-2
= (1/8) + (1/9) - (1/16)
lcm = 16*9
= 18/lcm + 16/lcm - 9/lcm
=25/(16*9)
= 25/144
====================
Cheers,
Stan H.